OpenStudy (anonymous):

please help i will fan and give a medal :] part a: Maria rented a coat at $285 for 3 days if she rents the same coat for 6 days she has to pay a total rent of $510 write an equation in the standard form to represent the total rent (y) that maria has to pay for renting the coat for x days Part b: write the equation obtained in part A using function notation part c: Describe the steps to graph the equation obtained above on the coordinate axes mention the lables on the axes and the intervals

OpenStudy (maddy1251):

I should be able to help.

OpenStudy (anonymous):

okay

OpenStudy (maddy1251):

What math level are you in?

OpenStudy (anonymous):

algebra 1

OpenStudy (maddy1251):

Sweet. Alright, I can help. :)

OpenStudy (anonymous):

okay thank you!

OpenStudy (maddy1251):

No need to thank me, I love math! :D

OpenStudy (maddy1251):

Let's just go step by step, we won't rush or jump to conclusions. So, part 1. part a: Maria rented a coat at $285 for 3 days if she rents the same coat for 6 days she has to pay a total rent of $510 write an equation in the standard form to represent the total rent (y) that maria has to pay for renting the coat for x days

OpenStudy (anonymous):

:D i do to though it depends on what type XD i am so bad at algebra

OpenStudy (maddy1251):

\[ax+bx=c\] This is standard form.

OpenStudy (anonymous):

okay

OpenStudy (maddy1251):

@snowblaze I worked really hard in Middle school to become good at math. Takes patience :)

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

so c is the cost

OpenStudy (maddy1251):

part a: Maria rented a coat at $285 for 3 days if she rents the same coat for 6 days she has to pay a total rent of $510 Okay so to get 285 for 3 days. That's 95 a day. For 510 is uhm I believe 85 a day if I did that mental math right. So I mean this gives us an idea of what the daily cost is, just to get numbers going. write an equation in the standard form to represent the total rent (y) that maria has to pay for renting the coat for x days

OpenStudy (anonymous):

86

OpenStudy (maddy1251):

Or, do you want to graph it.

OpenStudy (anonymous):

im personally better at equations

OpenStudy (maddy1251):

Okay

OpenStudy (maddy1251):

3x+b=285

OpenStudy (maddy1251):

6x+b=510

OpenStudy (anonymous):

so 510/6?

OpenStudy (maddy1251):

x=75

OpenStudy (anonymous):

okay

OpenStudy (maddy1251):

I think. Lemme double check that, too much mental math gets you in trouble.

OpenStudy (anonymous):

actually i think its 85

OpenStudy (maddy1251):

Well the slope is 75 between the two equations, now we have to move on to the next step.

OpenStudy (anonymous):

oh right sorry

OpenStudy (maddy1251):

\[y-y1=m(x-x1)\]

OpenStudy (maddy1251):

Please do not apologize :)

OpenStudy (maddy1251):

m=75

OpenStudy (anonymous):

:P okay

OpenStudy (maddy1251):

We can pick the points of (3, $285) to plug in x1=3 y1=285

OpenStudy (maddy1251):

soo\[y-285=75(x-3)\]

OpenStudy (anonymous):

okay

OpenStudy (maddy1251):

So they want it in y= I believe they said price for 'x' days. So we can solve so it is in y= form

OpenStudy (maddy1251):

y=75x+60

OpenStudy (maddy1251):

Part a is done : Maria rented a coat at $285 for 3 days if she rents the same coat for 6 days she has to pay a total rent of $510 y=75x+60

OpenStudy (anonymous):

okay

OpenStudy (maddy1251):

Maria rented a coat at $285 for 3 days if she rents the same coat for 6 days she has to pay a total rent of $510 Onto this.

OpenStudy (maddy1251):

Oops forgot the question in that. :P

OpenStudy (anonymous):

:P

OpenStudy (maddy1251):

It says to write in functional notation. that is simply saying 'f(x)'. Ever seen that before?

OpenStudy (anonymous):

yes

OpenStudy (maddy1251):

So replace 'y' with f(x) f(x)=75x+60

OpenStudy (anonymous):

ooooooh

OpenStudy (maddy1251):

Quite simple. :P haha

OpenStudy (maddy1251):

So now it's just c.

OpenStudy (anonymous):

okay

OpenStudy (maddy1251):

part c: Describe the steps to graph the equation obtained above on the coordinate axes mention the lables on the axes and the intervals Sweeeet. sounds fun. Do they really want you finding intervals of graphs?

OpenStudy (anonymous):

yurp

OpenStudy (maddy1251):

Well, I went ahead and graphed the function for you to visually see it.

OpenStudy (maddy1251):

I really do not see an interval..

OpenStudy (anonymous):

me neither

OpenStudy (maddy1251):

 Where I marked 'x' is considered intervals. This is a straight line.